Problem 10-25

Resistors for electronic circuits are manufactured on a high-speed automated machine. The machine is set up to produce a large run of resistors of 1,000 ohms each. Use Exhibit 10.13.

     To set up the machine and to create a control chart to be used throughout the run, 15 samples were taken with four resistors in each sample. The complete list of samples and their measured values are as follows: Use three-sigma control limits.

SAMPLE NUMBER READINGS (IN OHMS)
1 1097 1094 983 1026
2 978 1013 988 1014
3 976 991 1030 995
4 1016 991 1023 1014
5 977 1016 982 970
6 986 1011 1030 1027
7 1026 983 1017 1004
8 1017 1018 995 982
9 1010 981 994 1008
10 987 1020 1001 1004
11 994 995 1015 1026
12 1030 978 1017 984
13 974 995 1007 988
14 995 970 1004 1028
15 1012 1019 983 990

a. Calculate the mean and range for the above samples. (Round "Mean" to 2 decimal places and "Range" to the nearest whole number.)

Sample Number Mean Range
1
2
3
4
5
6
7
8
9
10
11
12
13
14
15

b. Determine X=X= and R−R−. (Round your answers to 3 decimal places.)

X=X=
R−R−

c. Determine the UCL and LCL for a X−X−chart. (Round your answers to 3 decimal places.)

UCL
LCL

d. Determine the UCL and LCL for R-chart. (Leave no cells blank - be certain to enter "0" wherever required. Round your answers to 3 decimal places.)

UCL
LCL
Answer #1
SAMPLE NUMBER READINGS (IN OHMS) X bar or Mean Minimum Maximum Range = Maximum - Minimum Upper Control Limit or UCL for X bar Chart Lower Control Limit or LCL for X bar Chart
1 1097 1094 983 1026 1050.00 983 1097 114 1037.601 971.699
2 978 1013 988 1014 998.25 978 1014 36 Upper Control Limit or UCL for R Chart Lower Control Limit or LCL for R Chart
3 976 991 1030 995 998.00 976 1030 54 103.146 0.000
4 1016 991 1023 1014 1011.00 991 1023 32 Here, 0.729, 2.282, 0 are constant and used here as Number of X = 4
5 977 1016 982 970 986.25 970 1016 46
6 986 1011 1030 1027 1013.50 986 1030 44
7 1026 983 1017 1004 1007.50 983 1026 43
8 1017 1018 995 982 1003.00 982 1018 36
9 1010 981 994 1008 998.25 981 1010 29
10 987 1020 1001 1004 1003.00 987 1020 33
11 994 995 1015 1026 1007.50 994 1026 32
12 1030 978 1017 984 1002.25 978 1030 52
13 974 995 1007 988 991.00 974 1007 33
14 995 970 1004 1028 999.25 970 1028 58
15 1012 1019 983 990 1001.00 983 1019 36
X Double Bar or Mean Value of X bar 1004.650 R bar or Mean Value of R 45.200

Formula

1 B C D E F G H I J K L
2 SAMPLE NUMBER READINGS (IN OHMS) X bar or Mean Minimum Maximum Range = Maximum - Minimum Upper Control Limit or UCL for X bar Chart Lower Control Limit or LCL for X bar Chart
3 1 1097 1094 983 1026 =AVERAGE(C3:F3) =MIN(C3:F3) =MAX(C3:F3) =I3-H3 =G18+(0.729*J18) =G18-(0.729*J18)
4 2 978 1013 988 1014 =AVERAGE(C4:F4) =MIN(C4:F4) =MAX(C4:F4) =I4-H4 Upper Control Limit or UCL for R Chart Lower Control Limit or LCL for R Chart
5 3 976 991 1030 995 =AVERAGE(C5:F5) =MIN(C5:F5) =MAX(C5:F5) =I5-H5 =2.282*J18 =0*J18
6 4 1016 991 1023 1014 =AVERAGE(C6:F6) =MIN(C6:F6) =MAX(C6:F6) =I6-H6 Here, 0.729, 2.282, 0 are constant and used here as Number of X = 4
7 5 977 1016 982 970 =AVERAGE(C7:F7) =MIN(C7:F7) =MAX(C7:F7) =I7-H7
8 6 986 1011 1030 1027 =AVERAGE(C8:F8) =MIN(C8:F8) =MAX(C8:F8) =I8-H8
9 7 1026 983 1017 1004 =AVERAGE(C9:F9) =MIN(C9:F9) =MAX(C9:F9) =I9-H9
10 8 1017 1018 995 982 =AVERAGE(C10:F10) =MIN(C10:F10) =MAX(C10:F10) =I10-H10
11 9 1010 981 994 1008 =AVERAGE(C11:F11) =MIN(C11:F11) =MAX(C11:F11) =I11-H11
12 10 987 1020 1001 1004 =AVERAGE(C12:F12) =MIN(C12:F12) =MAX(C12:F12) =I12-H12
13 11 994 995 1015 1026 =AVERAGE(C13:F13) =MIN(C13:F13) =MAX(C13:F13) =I13-H13
14 12 1030 978 1017 984 =AVERAGE(C14:F14) =MIN(C14:F14) =MAX(C14:F14) =I14-H14
15 13 974 995 1007 988 =AVERAGE(C15:F15) =MIN(C15:F15) =MAX(C15:F15) =I15-H15
16 14 995 970 1004 1028 =AVERAGE(C16:F16) =MIN(C16:F16) =MAX(C16:F16) =I16-H16
17 15 1012 1019 983 990 =AVERAGE(C17:F17) =MIN(C17:F17) =MAX(C17:F17) =I17-H17
18 X Double Bar or Mean Value of X bar =AVERAGE(G3:G17) R bar or Mean Value of R =AVERAGE(J3:J17)